If r, s, t are prime numbers and p, q are the positive integers such that their LCM of (p, q) is r^2 t^4 s^2 then the numbers of ordered pair of (p, q) is (A ) 252 ( B) 254 (C ) 225 (D ) 224

Grade:11

1 Answers

Vikas TU
8 years ago
exponent of r could be seleected as: => 1 + 2 + 2 = 5  ways.
For t = > 4 + 4 +1 = 9 ways.
for s = > 1 + 2 + 2 = 5 ways.
Therefore, 
no. of ordered pairs (p,q) choosen as: 5*9*5 = 225 ways.